# Framing global structural identifiability in terms of parameter symmetries

### v1

@@ 3.6.3 Generating and visualising parameter symmetries of the SEI model
First, we visualise the action of this symmetry on three parameter pairs (Fig. 1). These three parameter pairs are $(\mu_{E},\delta)$ for which the symmetry preserves the invariant $I_{3}$ in Eq. (99), $(k_{I},\beta)$ for which the symmetry preserves the invariant $I_{4}$ in Eq. (101) and $(\delta,\upsilon)$ for which the symmetry preserves the invariant $I_{5}$ in Eq. (103). Similarly, we visualise the action of this symmetry on two parameter triplets (Fig. 2). These triplets are $(\beta,\upsilon,c)$ for which the symmetry preserves the invariant $I_{6}$ in Eq. (106) and $(\beta,\delta,k_{E})$ for which the symmetry preserves the invariant $I_{7}$ in Eq. (109).
@@
Fig. 2: Three-dimensional projections of the action of a parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ of the SEI model. The action of the symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ generated by solving the system of ODEs in Eqs. (111)-(119) for the coefficients $(\alpha_{1},\alpha_{2})=(2,1)$ on six parameters illustrated by diamonds is visualised in two different three-dimensional subspaces as well as in two different two-dimensional subspaces of the nine-dimensional parameter space of the SEI model. These subspaces are (A) $(\beta,\upsilon,c)$ for which the invariant $I_{6}=\beta c\upsilon$ is preserved, (B) $(\beta,c\upsilon)$ for which the invariant $I_{6}=\beta c\upsilon$ is preserved, (C) $(\beta,\delta,k_{E})$ for which the invariant $I_{7}=\beta\delta/k_{E}$ is preserved and (D) $(\beta,\delta/k_{E})$ for which the invariant $I_{7}=\beta\delta/k_{E}$ is preserved.
@@ 4 Discussion
In this work, we have demonstrated how global structural identifiability can be understood in terms of the differential invariants of parameter symmetries. For the last two decades, the notion of classical Lie symmetries of ODEs acting on the independent and dependent variables by mapping solutions to other solutions [12, 13, 14, 15] has been extended to full symmetries which also account for rate parameters. Such full symmetries have been a large focus of research on the structural identifiability of mechanistic ODE models [6, 7, 8, 9, 10], and in particular a large emphasis has been put on developing algorithms for finding such symmetries in an automated fashion. However, the link between algebraic methods for global structural identifiability and symmetry based methods has, until this point, remained elusive. In this work, we established this conceptual link by introducing so-called parameter symmetries, Lie transformations that alter parameters while simultaneously preserving the observed outputs. In addition, we demonstrated that structural identifiability can be understood in terms of the differential invariants of these parameter symmetries. Based on these results, we proposed a three step recipe referred to as the CaLinInv-recipe which involves: (i) re-writing the original first-order ODE system as an equivalent ODE system for the outputs, also referred to as the Canonical coordinates; (ii) finding the parameter symmetries by solving the Linearised symmetry conditions; and (iii) elucidating the global structural identifiability by calculating the differential Invariants of the parameter symmetries. We later validated the CaLinInv-recipe by analysing the structural identifiability of two previously analysed mechanistic models of biological systems.
@@
The CaLinInv-recipe constitutes a new framing of the classical differential algebra approach for elucidating global structural identifiability in terms of Lie symmetries. The steps in this recipe are reminiscent of the differential algebra approach for global structural identifiability (Algorithm 1). In fact, the first steps in the differential algebra approach and the CaLinInv-recipe are identical, and this step attempts at finding algebraic equations relating inputs and outputs with rate parameters [2]. Technically, the differential algebra approach constructs a map between the rate parameters and the parameter combinations that can be inferred from the inputs and outputs, and then structural identifiability implies that this map is injective [21]. The parameter symmetries proposed in this work are essentially such maps, and the injectivity criterion can be understood in terms of the universal differential invariants of parameter symmetries. Better still, by framing global structural identifiability in terms of universal invariants of parameter symmetries, we understand why the standard differential algebra approach, which extracts coefficients in front of the monomials of the polynomial system of output ODEs, always finds identifiable parameter quantities, i.e. universal parameter invariants. This is due to the fact that the coefficients that are extracted in the differential algebra approach will either be a constant or a universal parameter invariant. This property is ensured by the definition of invariants of symmetries combined with the so-called linearised symmetry conditions, the equations defining parameter symmetries. In other words, the standard differential algebra approach is completely consistent with the notion of global structural identifiability expressed in terms of universal parameter invariants. Moreover, our symmetry-based approach for analysing global structural identifiability is theoretically generalisable to other mechanistic models consisting of, say, spatiotemporal systems of partial differential equations but in practice computer-assisted versions of this approach must be developed in order to analyse the global structural identifiability of such systems.
@@
An interesting future research direction is to automate the CaLinInv-recipe for systems of ODEs and eventually systems of partial differential equations. In the context of ODEs, it is known that if the right-hand sides or the reactions terms in the original first-order system are rational functions of the states, it is always possible to re-write the original system of first-order ODEs as a system of ODEs depending solely on the observed outputs [2]. Given such a re-formulated system in terms of the observed outputs, the two remaining steps of the CaLinInv recipe are straightforward to automate using symbolic calculations. This is also why the recipe can be automated, since many existing software for global structural identifiability, e.g. [22], conduct the first step of re-writing the original system so that it solely depends on the observed outputs in an automated fashion. Accordingly, the CaLinInv recipe can be implemented on top of existing algorithms for global structural identifiability analyses based on the differential algebra approach, which would result in an algorithm that not only yields the identifiable parameter combinations but also the family of parameter transformations that preserves the observed outputs, i.e. a family of parameter symmetries.
@@
In total, this work establishes a link between the existing body of work on full symmetries [6, 7, 8, 9, 10] and the differential algebra approach for global structural identifiability [2, 3, 4, 5]. Hitherto, it has been unclear how full symmetries transforming independent and dependent variables as well as parameters relate to global structural identifiability. The result which is closest to such a link was presented by Castro and de Boer [9] which states that a particular parameter is globally structurally identifiable if the only way to scale this parameter by a scaling factor that preserves the observed outputs, is if this scaling factor equals one. In fact, this is exactly what it means to say that the particular parameter of interest is a parameter invariant, and our theoretical framework based on parameter symmetries has formalised this result by demonstrating that a parameter is globally structurally identifiable if and only if it is a universal parameter invariant. Even better, our result generalises to any parameter symmetry as it is not restricted to the scalings studied by Castro and de Boer [9]. A succinct way of expressing our main result is that the globally structurally identifiable parameter quantities are given by universal parameter invariants. We have made a case for a perspective in which global structural identifiability is expressed in terms of differential invariants of parameter symmetries, and this work is a stepping stone towards fully exploiting the power of symmetry methods within the realm of global structural identifiability.
@@ Acknowledgements
JGB is funded by a grant from the Wenner-Gren foundations (Grant number: FT2023-0005). JBG thanks the Wenner-Gren foundations for a research fellowship. APB thanks the Mathematical Institute, Oxford for a Hooke Research Fellowship. This work was supported by a grant from the Simons Foundation (MP-SIP-00001828, REB)

### v3

@@ 4.5 Analysing local structural identifiability of an epidemiological model
First, we visualise the action of this symmetry on three parameter pairs (Fig. 1). These three parameter pairs are $(\mu_{E},\delta)$ for which the symmetry preserves the invariant $I_{3}$ in Eq. (80), $(k_{I},\beta)$ for which the symmetry preserves the invariant $I_{4}$ in Eq. (81) and $(\delta,\upsilon)$ for which the symmetry preserves the invariant $I_{5}$ in Eq. (82). Similarly, we visualise the action of this symmetry on two parameter triplets (Fig. 2). These triplets are $(\beta,\upsilon,c)$ for which the symmetry preserves the invariant $I_{6}$ in Eq. (83). and $(\beta,\delta,k_{E})$ for which the symmetry preserves the invariant $I_{7}$ in Eq. (84).
@@
Figure 2: Three-dimensional projections of the action of a parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ of the SEI model. The interpretation of these curves in parameter space is that the illustrated parameters have indistinguishable model behaviour. The action of the symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ generated by solving the system of ODEs in Eqs. (69)–(77) for the coefficients $(\alpha_{1},\alpha_{2})=(2,1)$ on six parameters illustrated by diamonds is visualised in two different three-dimensional subspaces as well as in two different two-dimensional subspaces of the nine-dimensional parameter space of the SEI model. These subspaces are (A) $(\beta,\upsilon,c)$ for which the invariant $I_{6}=\beta c\upsilon$ is preserved, (B) $(\beta,c\upsilon)$ for which the invariant $I_{6}=\beta c\upsilon$ is preserved, (C) $(\beta,\delta,k_{E})$ for which the invariant $I_{7}=\beta\delta/k_{E}$ is preserved and (D) $(\beta,\delta/k_{E})$ for which the invariant $I_{7}=\beta\delta/k_{E}$ is preserved.
@@ 5 Discussion
In this work, we have demonstrated how local structural identifiability can be understood in terms of the differential invariants of parameter symmetries. For the last two decades, the notion of classical Lie symmetries of ODEs acting on the independent and dependent variables by mapping solutions to other solutions [12, 13, 14, 15] has been extended to full symmetries which also account for parameters. Such full symmetries have been a large focus of research on the structural identifiability of mechanistic ODE models [6, 7, 8, 9, 10], and in particular a large emphasis has been put on developing algorithms for finding such symmetries in an automated fashion. However, the link between algebraic methods for determining structural identifiability and symmetry-based methods has, until this point, remained elusive. In this work, we establish this conceptual link by introducing so-called parameter symmetries, Lie transformations that alter parameters while simultaneously preserving the observed outputs. In addition, we demonstrate that local structural identifiability can be understood in terms of the differential invariants of these parameter symmetries. Based on these results, we propose a three step recipe referred to as the CaLinInv recipe which involves: (i) re-writing the original first-order ODE system as an equivalent ODE system for the outputs, also referred to as the Canonical coordinates; (ii) finding the parameter symmetries by solving the Linearised symmetry conditions; and (iii) elucidating the global structural identifiability by calculating the differential Invariants of the parameter symmetries. We later demonstrated practical use of the CaLinInv recipe by analysing the local structural identifiability of two previously analysed mechanistic models of biological systems.
@@
The CaLinInv recipe constitutes a new framing of the classical differential algebra approach for elucidating structural identifiability in terms of Lie symmetries. The steps in this recipe are reminiscent of the differential algebra approach for structural identifiability. In fact, the first steps—finding algebraic equations relating inputs and outputs with parameters [2]—are identical. Technically, the differential algebra approach constructs a map between the parameters and the parameter combinations that can be inferred from the inputs and outputs, and then structural identifiability requires that this map is injective [23]. The parameter symmetries proposed in this work are a way of constructing such maps, and the injectivity criterion can be understood in terms of the universal differential invariants of parameter symmetries. By framing structural identifiability in terms of universal invariants of parameter symmetries, we understand why the standard differential algebra approach, which extracts coefficients in front of the monomials of the polynomial system of output ODEs, always finds identifiable parameter quantities; that is, universal parameter invariants. This is due to the fact that the coefficients that are extracted in the differential algebra approach will either be a constant or a universal parameter invariant. This property is ensured by the definition of invariants of symmetries combined with the so-called linearised symmetry conditions, the equations defining parameter symmetries. In other words, the standard differential algebra approach is consistent with the notion of local structural identifiability expressed in terms of universal parameter invariants. Moreover, our symmetry-based approach for analysing local structural identifiability is theoretically generalisable to other mechanistic models consisting of, say, systems of partial differential equations. However, in practice, computer-assisted versions of this approach must be developed in order to analyse the structural identifiability of such systems. In particular, two aspects of the symmetry-based structural identifiability analysis should be automated, namely the formulation of the input-output system and an implementation for solving the linearised symmetry conditions.
@@
An interesting future research direction is to automate the CaLinInv recipe for systems of ODEs and eventually systems of partial differential equations. In the context of ODEs, it is known that if the right-hand sides of the original first-order system are rational functions of the states it is always possible to re-write the original system of first-order ODEs as a system of ODEs depending solely on the observed outputs [2]. Given such a re-formulated system in terms of the observed outputs, the two remaining steps of the CaLinInv recipe are straightforward to automate using symbolic calculations. This is also why the recipe can be automated, since many existing software for structural identifiability (e.g. [24]) conduct the first step of re-writing the original system so that it solely depends on the observed outputs in an automated fashion. Accordingly, the CaLinInv recipe can be implemented on top of existing algorithms for structural identifiability analyses based on the differential algebra approach, which would result in an algorithm that not only yields the identifiable parameter combinations but also the family of parameter transformations that preserves the observed outputs; that is, a family of parameter symmetries.
@@
Altogether, this work establishes a link between the existing body of work on full symmetries [6, 7, 8, 9, 10] and the differential algebra approach for structural identifiability [2, 3, 4, 5]. Until now, it has been unclear how the full symmetries transforming independent and dependent variables as well as parameters relate to structural identifiability. The result which is closest to such a link was presented by Castro and de Boer [9] which states that a particular parameter is globally structurally identifiable if the only way to scale this parameter by a scaling factor that preserves the observed outputs, is if this scaling factor equals one. In fact, this is exactly what it means to say that the particular parameter of interest is a parameter invariant, and our theoretical framework based on parameter symmetries has formalised this result by demonstrating that a parameter is locally structurally identifiable if and only if it is a universal parameter invariant. Indeed, our result generalises to any parameter symmetry as it is not restricted to the scalings studied by Castro and de Boer [9]. As has been noted previously, by only focusing on scalings one misses other important parameter transformations that preserve observed outputs [11], and thus the conclusions about structural identifiability drawn from studying such scalings can be misleading. In addition, Lie symmetries acting on states and parameters have previously been calculated in an automated fashion [8], but the exact link between structural identifiability and these calculated symmetries has been unclear. A succinct way of expressing this link in words which is our main result is that the locally structurally identifiable parameter quantities are given by universal parameter invariants. We have made a case for a perspective in which local structural identifiability is expressed in terms of differential invariants of parameter symmetries, and this work is a stepping stone towards fully exploiting the power of symmetry methods within the realm of local structural identifiability.
@@ Acknowledgements
JGB is funded by a grant from the Wenner-Gren foundations (Grant number: FT2023-0005). JBG thanks the Wenner-Gren foundations for a research fellowship. APB thanks the Mathematical Institute, Oxford for a Hooke Research Fellowship. This work was supported by a grant from the Simons Foundation (MP-structural identifiabilityP-00001828, REB). or the purpose of open access, the author has applied a CC BY public copyright licence to any author accepted manuscript arising from this submission.